OpenStudy (anonymous):

a kayaker paddles upstream from camp to photograph a waterfall and returns.the kayakers speed while traveling upstream and downstream is shown below.what is the kayakers speed in still water?what is the speed of the current????????? HELPP

OpenStudy (anonymous):

is there supposed to be a diagram? O.o

OpenStudy (anonymous):

yea there is but it wont let me put it

OpenStudy (anonymous):

if u search up the question a pdf file should pop up and then u just click on it and the picture is there

OpenStudy (anonymous):

this one?

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

give me a sec. i'll try to solve it on my own

OpenStudy (anonymous):

okay thankss

OpenStudy (anonymous):

oh, this is linear systems?! lol, i thought this is vectors -.- okay.. so let x be the current's speed and y be the boat's speed in still water the current's direction is downstream first let's do an equation for upstream since the kayaker went upstream at a speed of 4mi/h and the kayaker's direction is opposite to the current's direction, we will have an equation of \(x-y=4\) this will be the first equation next is the equation for downstream kayaker's speed=7mi/h have same direction with the current \(x+y=7\) this will be our second equation so we have \(x-y=4\) \(x+y=7\) now use elimination or substitution to solve for x and y :)

OpenStudy (anonymous):

oh okay tysm :))

OpenStudy (anonymous):

did you get it? ^.^

OpenStudy (anonymous):

how do i use substition or elimination for that ?

OpenStudy (anonymous):

elimination is easier you need to eliminate one variable to solve for the other one for example you want to solve for "x" you just add the two equations and the variable y will be eliminated

OpenStudy (anonymous):

ohohoh so i just add 4+7=11?

OpenStudy (anonymous):

ohhh i get it

OpenStudy (anonymous):

err.. not really new equation will be : 2x=11 then just solve for x after you solve for the value of x, substitute it in any of the two equations

OpenStudy (anonymous):

it would be 2x=11 and then divide that which gives me 5.5

OpenStudy (anonymous):

omg tysm

OpenStudy (anonymous):

so thats it right

OpenStudy (anonymous):

yes, then you need to solve for "y"

OpenStudy (anonymous):

so i do 5.5-y=4

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh ok omg ur so smart

OpenStudy (anonymous):

i may need ur help for other excersices if thats okay with u

OpenStudy (anonymous):

lol, you already get this one? O.o

OpenStudy (anonymous):

yes i do! :) but i get confused really easily lol

OpenStudy (anonymous):

oh ok the answer is 5.5 mi/h;1.5 mi/h

OpenStudy (anonymous):

that's fine, you'll get used to it if you'll practice a lot also, you may get a negative speed sometimes, which only refers to the direction of the object

OpenStudy (anonymous):

ugh i just hate algebra but thanks alot u saved me and oh ok :)

OpenStudy (anonymous):

no problem :)